# Evolution of the Inner Accretion Flow in Swift J1727.8–1613 across Intermediate States: Insights from Broadband Spectral and Timing Analysis

### v1

@@ 2.1 NICER
We processed the NICER (Gendreau2016SPIE.9905E..1HG) observations using the standard task nicerl2 available within HeaSoft (v.6.34) and the most recent CALDB (v.xti_20240206). To extract the time-averaged source and background energy spectra, we used the standard script nicerl3-spec, where the background spectra are estimated using the ‘3C50’ model. We further used the nicerl3-lc task to generate the lightcurves from both the observations.
2.2 NuSTAR
@@
The NuSTAR (Harrison2013ApJ...770..103H) data were processed using the standard task nupipeline, which is part of the NuSTARDAS software package, along with the latest calibration files. Since the source was very bright during these observations, we applied the status expression statusexpr=‘‘(STATUS==b0000xxx00xxxx000)&&     (SHIELD==0)’’ during data processing. Source and background spectra were then extracted considering circular regions of $70\arcsec$ radius, centered at the source position and an off-source region, respectively, using the standard task nuproducts. The corresponding response matrix files (RMFs) and auxiliary response files (ARFs) were generated using the same task. It is worth noting that the NuSTAR detectors may be affected by dead time [^1] It is the interval after each detected event during which the instrument processes the signal and cannot record new photons., particularly for bright sources. While dead time does not have any effect on spectroscopy, it can significantly impact timing studies (Bachetti2015ApJ...800..109B; Bachetti2018ApJ...853L..21B, for details on NuSTAR dead time).
2.3 AstroSat/CZTI
@@
AstroSat, India’s first multi-wavelength astronomy satellite, was launched on September 28, 2015. It carries four co-aligned instruments onboard: the Soft X-ray Telescope (SXT; Singh2016SPIE.9905E..1ES; Singh2017JApA...38...29S), the Large Area X-ray Proportional Counter Array (LAXPC; Yadav2016SPIE.9905E..1DY; yadav2016ApJ...833...27Y; Agrawal2017JApA...38...30A; Antia2017ApJS..231...10A), Cadmium-Zinc-Telluride Imager (CZTI) (Vadawale2016SPIE.9905E..1GV; Vibhute2021JApA...42...76V), and the Ultra-Violet Imaging Telescope(UVIT; Tandon2017; Tandon2020). For this study, we utilized data exclusively from the CZTI instrument. We downloaded the CZTI data from the AstroSat data archive [^2] https://astrobrowse.issdc.gov.in/astro_archive/archive/Home.jsp and produced clean event files using the latest version of the cztpipeline (v3.0) [^3] http://astrosat-ssc.iucaa.in/cztiData and updated calibrations. This version of the cztpipeline includes an improved mask-weighting technique implemented in the cztbindata module (Chattopadhyay2024ApJ...960L...2C). The clean event files were then used to extract background-subtracted spectra for each quadrant using the standard tasks within cztpipeline [^4] http://astrosat-ssc.iucaa.in/uploads/threadsPageNew_SXT.html. Finally, spectra from all four quadrants were combined to produce a high signal-to-noise ratio spectrum. We then added a $2\%$ systematic error to the CZTI merged spectral data to account for the calibration uncertainties of the instrument. We analyzed the CZTI spectral data in the 4040200200 keV band, excluding data below $40$ keV due to the strong presence of Xe hump, which peaks around $\sim 30$ keV.
@@ 2.4 Insight-HXMT/HE
The first X-ray satellite of China, Insight-HXMT (Zhang2014SPIE.9144E..21Z; Zhang2018SPIE10699E..1UZ; Zhang2020SCPMA..6349502Z), launched on June 15, 2017, carries three different payloads onboard, and they are: Low-Energy X-ray Telescope (LE; Chen2020SCPMA..6349505C), Medium-Energy X-ray Telescope (ME; Cao2020SCPMA..6349504C), and High-Energy X-ray Telescope (HE; Liu2020SCPMA..6349503L). Here, we have utilized the data only from HE, which operates in the 2025020-250 keV band, for spectral analysis. We used HXMTDAS (v.2.06; Zhao2020ASPC..527..469Z) package to process the HE data following the standard procedure and default inputs. For the generation of source and background spectra and the response files, the standard task hpipeline has been used. For this study, we utilized the HE spectral data in the 3030150150 keV band and discarded the data below $30$ keV due to calibration issues, and beyond $150$ keV for significant background contamination.
3 Analysis and Results
3.1 Broadband Spectral Analysis
@@
We performed broadband spectral analysis using XSPEC (v.12.14.1; Arnaud1996ASPC..101...17A) and quoted all the uncertainties on the best-fit parameters at the $90\%$ confidence ($\Delta\chi^{2}\approx 2.71$) level. We used the optimal binning algorithm (Kaastra2016A&A...587A.151K) with a minimum of 25 counts per grouped bin to the NICER, NuSTAR, and HE spectral data to enable the use of $\chi^{2}$ statistics. For the CZTI spectral data, we retained the default binning criteria embedded in the cztpipeline (Chattopadhyay2024ApJ...960L...2C; Chand2024ApJ...972...20C). To account for any differences in the relative normalizations and possible calibration uncertainties between the different instruments, we multiplied a plabs component to our models. This component is purely phenomenological and does not alter the intrinsic shape of the continuum or the key physical parameters. The parameters $\Delta\Gamma$ and $K$ of the plabs component were fixed at 0 and $1$, respectively, for NuSTAR FPMA, while they were varied freely for the other instruments. We started the spectral analysis by jointly fitting the simultaneous (for Obs1) or contemporaneous (for Obs2) NICER and NuSTAR spectral data in the 0.7780.7-78 keV band with a model consisting of a multicolor disk blackbody component (diskbb; Mitsuda1984PASJ...36..741M) to account for the thermal emission from the disk, and thcomp (Zdziarski2020MNRAS.492.5234Z), describing the spectra from Comptonization of the disk photons by thermal electrons in the corona. As thcomp is a convolution model, we extended the sample energy range from $0.01$ to $1000$ keV using the energies command within XSPEC. Furthermore, we treated the absorption by the interstellar medium using the Galactic absorption component tbabs with the abundances and cross sections taken from Wilms2000ApJ...542..914W and Verner1996ApJ...465..487V, respectively. This model plabs*tbabs(thcomp*diskbb) provided a very poor spectral fit and left prominent residuals at the Fe $K_{\alpha}$ band and reflection hump region, typically arising due to the reflection of the coronal irradiation from the accretion disk. Figure 3 shows the NuSTAR/FPMA spectral data for the two different spectral states along with the data residuals to the model tbabs(thcomp*diskbb). A clear difference in the spectral nature between the two spectral states can be seen from this figure. The nature of the reflection hump also shows a transition over the two observations, where it appears to be relatively prominent at high energies in the Obs1 and weak in the Obs2.
@@
Figure 3: Upper panel: NuSTAR/FPMA spectral data in the HIMS (blue), the SIMS (red), and the fitted model tbabs*(thcomp*diskbb) (black solid line). Middle panel: data residuals with respect to the fitted model. Lower panel: data-to-model ratio. The presence of a prominent Fe $K_{\alpha}$ line and a weak reflection hump is seen.
3.1.1 HIMS
@@
$R_{\rm{in}}$ ($r_{\mathrm{g}}$) $143.7_{-110.4}$  $\cdots$  $\cdots$
@@
$A_{\mathrm{Fe}}$ (solar) $0.61^{+0.44}$  $\cdots$  $\cdots$
$\log N$ $15.98^{+0.51}$  $\cdots$  $\cdots$
@@
mbknpo $B$ $1.0^{+0.12}$  $\cdots$  $\cdots$
@@
Relconv $R_{\rm{in}}$ ($r_{\mathrm{g}}$)  $\cdots$ $249.5_{-196.4}$ $212.6_{-168.8}$
@@
$\Delta\Gamma_{\rm{NuSTAR~B}}$ $004^{+0.005} 0.0$$04^{+0.005} 0.0$$004^{+0.005}$
@@
  • Notes – M1: plabs*tbabs(thcomp*diskbb+mbknpo*relxillCp), M2: plabs*tbabs*relconv*xilconv(thcomp*diskbb+compps[thermal]), M3: plabs*tbabs*relconv*xilconv*(thcomp*diskbb+compps[hybrid]), where the spectral data from HE and CZTI were also added. See Svoboda2024ApJ...966L..35S for the definition of the mbknpo in xspec. The spin of the black hole and disk inclination angle are fixed at $0.998$ and $40^{\circ}$, respectively. f – indicates fixed parameters.
@@
For Obs1, we then attempted to model reflection from ionized material by convolving ireflect with thcomp and incorporating a gaussian component to account for the Fe $K_{\alpha}$ emission line. The ireflect model is the generalized convolution form of pexriv, and describes only the hard X-ray shape of the reflected spectrum, excluding all the line emissions. However, this model configuration, plabs*tbabs(ireflect*thcomp*diskbb+gaussian), yielded a statistically poor fit, with $\chi^{2}/dof=1634.4/596$, and left significant residuals at high energies. This indicates that a model involving a single Comptonizing region is insufficient to reproduce the observed broadband spectral shape, and an additional Comptonization component may be required to account for the high-energy excess. A requirement for two distinct Comptonizing regions to describe the broadband (0.71000.7-100 keV) spectrum in the LHS of the low-mass BHXRBs GX 339–4 (Chand2024ApJ...972...20C) and AT2019wey (Sahu2025arXiv251106393S) has also been reported by our group . A similar behavior is observed in the LHS of MAXI J1820+070, indicating that multiple Comptonization regions are required in the hard states of BHXRBs (2021ApJ...909L...9Z; Zdziarski2021ApJ...914L...5Z).
@@
We then removed ireflect and gaussian components and adopted a more physically motivated model relxillCp, which is a part of the family of relativistic blurred reflection models relxill (Dauser2014MNRAS. [^4] L.100D; Garcia2014ApJ...782...76G). The relxillCp model self-consistently computes the reflection spectrum arising from the irradiation of thermally Comptonized coronal emission from the disk and also includes relativistic effects due to strong gravity near the black hole. The shape of the incident radiation in relxillCp is the same as in the nthcomp model. In relxillCp, we fixed the black hole spin at the maximum value of $0.998$ and disk inclination angle at $40^{\circ}$ (Peng2024ApJ...960L..17P; Svoboda2024ApJ...966L..35S). The outer disk radius was fixed at $1000~r_{\rm{g}}$, and a single emissivity profile ($\propto r^{-q}$, $q$ being the emissivity index) was considered throughout the whole disk with $q$ fixed at the canonical value of 3. The photon index ($\Gamma$) and electron temperature ($kT_{\rm{e}}$) were varied independently of those in thcomp. Other free parameters of relxillCp were the inner disk radius ($R_{\rm{in}}$), log of disk ionization ($\log\xi$), iron abundance ($A_{\rm{Fe}}$), log of disk density ($\log N$), reflection fraction ($R_{\rm{ref}}$), and normalization. To avoid the over prediction of the contribution of the reflection at lower energies, we multiplied relxillCp with a mbknpo component, which helps to put a break on the reflection spectrum below a certain energy, and thus preventing the unphysical runaway of the photon spectrum at low energies (Steiner2024ApJ...969L..30S; Svoboda2024ApJ...966L..35S). The mbknpo component is parameterized with the break energy $B$ and index $I$, used for the correction for energies $E<B$. We set these parameters following the prescription outlined by Svoboda2024ApJ...966L..35S. This composite model, plabs*tbabs(thcomp*diskbb+mbknpo*relxillCp) (hereafter M1), resulted in a significantly improved fit with $\chi^{2}/dof=457.3/590$. The best-fit spectral parameters are listed in Table 2, and the corresponding spectral models and data are shown in Figure 4.
@@
The results from our model M1 clearly indicate that the broadband spectrum in the 0.7780.7-78 keV band in the HIMS requires two separate Comptonizing regions. The relxillCp component represents the hard Comptonizing region with $\Gamma\sim 1.8$ and $kT_{\rm{e}}\sim 16$ keV, likely having a spherically extended geometry. This region is primarily responsible for the origin of the reflection features at a distant part of the disk. On the other hand, the second Comptonizing region, described by thcomp, is found to have low electron temperature, $kT_{\rm{e}}\sim 4$ keV, and exhibits a relatively steeper spectrum with $\Gamma\sim 1.95$. This soft Comptonizing region may appear as a warm layer over the inner parts of the accretion disk. We also explored the possibility of an additional contribution to the reflection from the soft Comptonizing region by incorporating a second relxillCp as a reflection component only. However, we did not find any significant improvement in our spectral fitting. This implies that the reflection associated with the soft Comptonizing region is negligible, and the overall reflection spectrum is dominantly governed by the hard Comptonizing region.
@@
The above model could not constrain the inner disk radius, $R_{\rm{in}}$, properly. However, the lower bound at the $90\%$ confidence level indicates that the disk is truncated away from the ISCO by at least $\sim 33~r_{\rm{g}}$ (where $r_{\rm{g}}=GM/c^{2}$ is the Gravitational radius). We also calculated the inner disk radius from the best-fit value of the diskbb normalization following its definition in XSPEC and considering the color correction factor of 1.7 (Shimura1995ApJ...445..780S). The estimated radius is found to be $\sim 29~r_{\rm{g}}$, which agrees closely with the obtained value from the reflection modeling. Furthermore, the temperature of the inner accretion disk is found to be $\sim 0.4$ keV, which is typically higher than that observed in the LHS. The disk appears to be highly ionized with slightly elevated density. Within the uncertainties, the iron abundance is consistent with the solar value. Additionally, the reflection fraction, $R_{\rm{ref}}$, defined as the ratio of the coronal emission reaching the disk and the observer, is found to be $\sim 0.64$, which in turn also favors the disk truncation scenario. All the best-fit spectral parameters indicate that the source was in HIMS during Obs1.
@@
We then modeled the hard Comptonization spectra with the compps (Poutanen1996ApJ...470..249P) model by replacing relxillCp in M1. The Comptonization model compps can treat the electron distribution in a corona as purely thermal, purely non-thermal, or hybrid for different shapes of the coronal geometry. Here, we assumed the electron distribution in the corona as purely thermal and the coronal geometry as spherical (geom=0) in the compps model. We also tested other geometries of the corona but did not notice any significant change in the spectral fitting. Therefore, we kept the geometry spherical, which also provides a relatively faster method for fitting the spectra than with the other geometries. In the compps, the seed photons for the Comptonization are considered to be multi-color disk photons provided by the diskbb component. We fixed the disk inclination angle at $40^{\circ}$ and iron abundance at solar. The other free parameters of this model component are the electron temperature, $kT_{\rm{e}}$, and the optical depth, $\tau$, of the scattering medium. Furthermore, to account for the reflection from the disk, we convolved the xilconv model with both the Comptonizing regions. The reflection fraction, $R_{\rm{refl}}$, and log of disk ionization, $\log\xi$, parameters in compps were allowed to vary freely. We kept the iron abundance fixed at solar, consistent with the value obtained from the model M1 within the uncertainty level. We also incorporated the relconv model to treat the relativistic blurring effects, while keeping the relevant parameters the same as those used for relxillCp in model M1. This model, plabs*tbabs*relconv*xilconv*(thcomp*diskbb+compps) (referred to as M2), provides a statistically acceptable fit to the broadband spectrum, with $\chi^{2}/dof=458.7/592$. The best-fit spectral parameters are given in Table 2, and the best-fit spectra along with model components are shown in Figure 4.
@@
In the next scenario, we incorporated spectral data from the HE and CZTI instruments onboard Insight-HXMT and AstroSat, respectively, to achieve broadband spectral coverage up to 200 keV. We first checked for any calibration uncertainties present in these data sets by fitting them up to $78$ keV along with NICER and NuSTAR spectra data with both models M1 and M2. The resulting best-fit spectral parameters were either the same or well consistent with those found from the joint spectral fitting of NICER and NuSTAR spectral data in the 0.7780.7-78 keV band using models M1 and M2, indicating no significant cross-calibration discrepancies among the instruments. Furthermore, inspection of the HE spectral data confirms that the background is significantly lower than the source rate in the 100150100-150 keV band. Based on the source and background count rates and their statistical uncertainties, we estimate a detection significance of $\sim 52\sigma$ for the source emission in the 100150100-150 keV band. We then extended the energy ranges of the HE and CZTI spectral data to $150$ keV and $200$ keV, respectively. However, we find that both models M1 and M2 were unable to describe the broadband spectra up to $200$ keV, and provided poor spectral fits with $\chi^{2}$/dof=803/679=803/679 for M1 and $886.8/682$ for M2, leaving a prominent hump at high energies ($\gtrsim 100$ keV). Figure 5 presents the broadband spectral data in the 0.72000.7-200 keV band fitted with the model M2, along with the residuals, highlighting the deviations between the observed data and the model.

### v2

@@ II.1 NICER
We processed the NICER (Gendreau et al., 2016) observations using the standard task nicerl2 available within HeaSoft (v.6.34) and the most recent CALDB (v.xti_20240206). To extract the time-averaged source and background energy spectra, we used the standard script nicerl3-spec, where the background spectra are estimated using the ‘3C50’ model. We further used the nicerl3-lc task to generate the lightcurves from both the observations.
II.2 NuSTAR
@@
The NuSTAR (Harrison et al., 2013) data were processed using the standard task nupipeline, which is part of the NuSTARDAS software package, along with the latest calibration files. Since the source was very bright during these observations, we applied the status expression statusexpr=‘‘(STATUS==b0000xxx00xxxx000)&&     (SHIELD==0)’’ during data processing. Source and background spectra were then extracted considering circular regions of $70\arcsec$ radius, centered at the source position and an off-source region, respectively, using the standard task nuproducts. The corresponding response matrix files (RMFs) and auxiliary response files (ARFs) were generated using the same task. It is worth noting that the NuSTAR detectors may be affected by dead time [^1] It is the interval after each detected event during which the instrument processes the signal and cannot record new photons., particularly for bright sources. While dead time does not have any effect on spectroscopy, it can significantly impact timing studies (Bachetti et al., 2015; Bachetti and Huppenkothen, 2018, for details on NuSTAR dead time).
II.3 AstroSat/CZTI
@@
AstroSat, India’s first multi-wavelength astronomy satellite, was launched on September 28, 2015. It carries four co-aligned instruments onboard: the Soft X-ray Telescope (SXT; Singh et al., 2016, 2017), the Large Area X-ray Proportional Counter Array (LAXPC; Yadav et al., 2016a, b; Agrawal et al., 2017; Antia et al., 2017), Cadmium-Zinc-Telluride Imager (CZTI) (Vadawale et al., 2016; Vibhute et al., 2021), and the Ultra-Violet Imaging Telescope(UVIT; Tandon et al., 2017, 2020). For this study, we utilized data exclusively from the CZTI instrument. We downloaded the CZTI data from the AstroSat data archive [^2] https://astrobrowse.issdc.gov.in/astro_archive/archive/Home.jsp and produced clean event files using the latest version of the cztpipeline (v3.0) [^3] http://astrosat-ssc.iucaa.in/cztiData and updated calibrations. This version of the cztpipeline includes an improved mask-weighting technique implemented in the cztbindata module (Chattopadhyay et al., 2024). The clean event files were then used to extract background-subtracted spectra for each quadrant using the standard tasks within cztpipeline [^4] http://astrosat-ssc.iucaa.in/uploads/threadsPageNew_SXT.html with the default inputs. For CZTI, background subtraction is inherent to spectral extraction and is based on analysis of long blank-sky observations and the mask-weighting technique. For typical exposures of less than $100$ ks, the background subtraction systematics are negligible. This becomes comparable to the statistical uncertainties only for exposures of $500$ ks and higher [^5] http://astrosat-ssc.iucaa.in/uploads/meetings/13.%20CZTISpecCalStatus_Mithun%20-%20Mithun%20N%20P%20S.pdf (N. P. S. Mithun from the CZTI team, private communication). Finally, spectra from all four quadrants were combined to produce a high-signal-to-noise spectrum. We then added a 2% systematic error to the CZTI merged spectral data to account for the instrument’s calibration uncertainties. We analyzed the CZTI spectral data in the 40–200 keV band, excluding data below 40 keV because of the strong Xe hump, which peaks at $\sim 30$ keV.
@@ II.4 Insight-HXMT/HE
The first X-ray satellite of China, Insight-HXMT (Zhang et al., 2014, 2018, 2020), launched on June 15, 2017, carries three different payloads onboard. They are: Low-Energy X-ray Telescope (LE; Chen et al., 2020), Medium-Energy X-ray Telescope (ME; Cao et al., 2020), and High-Energy X-ray Telescope (HE; Liu et al., 2020). Here, we have used only HE data, which operates in the 20–250 keV band, for spectral analysis. We used HXMTDAS (v.2.06; Zhao, 2020) package to process the HE data following the standard procedure and default inputs. For generating source and background spectra and response files, the standard task hpipeline was used. The background is generated and scaled by the standard task hebkgmap within hpipeline, which employs the built-in background model by Liao et al. (2020) based on blank-sky observations. This background model introduces an average systematic error of 1.5% in the background energy spectrum (26–-100 keV) for a typical exposure of 8 ks. The background systematic error varies in the range of 2%–10% above 120 keV (Li et al., 2020). For this study, we utilized the HE spectral data in the 30–150 keV band and discarded the data below $30$ keV due to calibration issues, and beyond 150 keV for significant background contamination.
III Analysis and Results
III.1 Broadband Spectral Analysis
@@
We performed broadband spectral analysis using XSPEC (v.12.14.1; Arnaud, 1996) and quoted all the uncertainties on the best-fit parameters at the $90\%$ confidence ($\Delta\chi^{2}\approx 2.71$) level. We applied the optimal binning algorithm (Kaastra and Bleeker, 2016), with a minimum of 25 counts per bin, to the NICER, NuSTAR, and HE spectral data to enable the use of the $\chi^{2}$ statistic. For the CZTI spectral data, we retained the default binning criteria embedded in the cztpipeline (Chattopadhyay et al., 2024; Chand et al., 2024). To account for any differences in the relative normalizations and possible calibration uncertainties between the different instruments, we multiplied a plabs component to our models. This component is purely phenomenological and does not alter the continuum’s intrinsic shape or the key physical parameters. The parameters $\Delta\Gamma$ and $K$ of the plabs component were fixed at 0 and $1$, respectively, for NuSTAR FPMA, while they were varied freely for the other instruments. We started the spectral analysis by jointly fitting the simultaneous (for Obs1) or contemporaneous (for Obs2) NICER and NuSTAR spectral data in the 0.7780.7-78 keV band with a model consisting of a multicolor disk blackbody component (diskbb; Mitsuda et al., 1984) to account for the thermal emission from the disk, and thcomp (Zdziarski et al., 2020), describing the spectra from Comptonization of the disk photons by thermal electrons in the corona. As thcomp is a convolution model, we extended the sample energy range from 0.01 to 1000 keV using the energies command in XSPEC. Furthermore, we treated the absorption by the interstellar medium using the Galactic absorption component tbabs with the abundances and cross sections taken from Wilms et al. (2000) and Verner et al. (1996), respectively. This model plabs*tbabs(thcomp*diskbb) provided an inferior spectral fit and left prominent residuals at the Fe $K_{\alpha}$ band and reflection hump region, typically arising due to the reflection of the coronal irradiation from the accretion disk. Figure 3 shows the NuSTAR/FPMA spectral data for the two different spectral states along with the data residuals to the model tbabs(thcomp*diskbb). A clear difference in spectral properties between the two states is evident in this figure. The nature of the reflection hump also shows a transition between the two observations, with it appearing relatively prominent at high energies in Obs1 and relatively weak in Obs2.
@@
Figure 3: Upper panel: NuSTAR/FPMA spectral data in the HIMS (blue), the SIMS (red), and the fitted model tbabs*(thcomp*diskbb) (black solid line). Middle panel: data residuals with respect to the fitted model. Lower panel: data-to-model ratio. A prominent Fe $K_{\alpha}$ line and a weak reflection hump are observed.
III.1.1 HIMS
@@
$R_{\rm{in}}$ ($r_{\mathrm{g}}$) $143.7^{p}_{-110.4}$  $\cdots$  $\cdots$
@@
$A_{\mathrm{Fe}}$ (solar) $0.61^{+0.44}_{p}$  $\cdots$  $\cdots$
$\log N$ $15.98^{+0.51}_{p}$  $\cdots$  $\cdots$
@@
mbknpo $B$ $1.0^{+0.12}_{p}$  $\cdots$  $\cdots$
@@
Relconv $R_{\rm{in}}$ ($r_{\mathrm{g}}$)  $\cdots$ $249.5^{p}_{-196.4}$ $212.6^{p}_{-168.8}$
@@
$\Delta\Gamma_{\rm{NuSTAR~B}}$ $004^{+0.005}_{p} 0.0$$04^{+0.005}_{p} 0.0$$004^{+0.005}_{p}$
@@
  • Notes – M1: plabs*tbabs(thcomp*diskbb+mbknpo*relxillCp), M2: plabs*tbabs*relconv*xilconv(thcomp*diskbb+compps[thermal]), M3: plabs*tbabs*relconv*xilconv*(thcomp*diskbb+compps[hybrid]), where the spectral data from HE and CZTI were also added. See Svoboda et al. (2024) for the definition of the mbknpo in xspec. The spin of the black hole and disk inclination angle are fixed at $0.998$ and $40\arcdeg$, respectively. f – indicates fixed parameters. $p$ – indicates that the parameter is pegged at the model-defined highest or lowest bound during error estimation.
@@
For Obs1, we then attempted to model reflection from ionized material by convolving ireflect (Magdziarz and Zdziarski, 1995) with thcomp and incorporating a gaussian component to account for the Fe $K_{\alpha}$ emission line. The ireflect model is the generalized convolution form of pexriv, and describes only the hard X-ray shape of the reflected spectrum, excluding all the line emissions. However, this model configuration, plabs*tbabs(ireflect*thcomp*diskbb+gaussian), yielded a statistically poor fit, with $\chi^{2}/dof=1634.4/596$, and left significant residuals at high energies. This indicates that a model involving a single Comptonizing region is insufficient to reproduce the observed broadband spectral shape, and an additional Comptonization component may be required to account for the high-energy excess. A requirement for two distinct Comptonizing regions to describe the broadband (0.71000.7-100 keV) spectrum in the LHS of the low-mass BHXRBs GX 339–4 (Chand et al., 2024) and AT2019wey (Sahu et al., 2025) has also been reported by our group. A similar behavior is observed in the LHS of MAXI J1820+070, indicating that multiple Comptonization regions are required in the hard states of BHXRBs (Zdziarski et al., 2021a, c).
@@
We then removed ireflect and gaussian components and adopted a more physically motivated model relxillCp, which is a part of the family of relativistic blurred reflection models relxill (Dauser et al., 2014; García et al., 2014). The relxillCp model self-consistently computes the reflection spectrum arising from the irradiation of thermally Comptonized coronal emission from the disk and also includes relativistic effects due to strong gravity near the black hole. The shape of the incident radiation in relxillCp is the same as in the nthcomp model. In relxillCp, we fixed the black hole spin at the maximum value of $0.998$ and disk inclination angle at $40\arcdeg$ (Peng et al., 2024; Svoboda et al., 2024). The outer disk radius was fixed at $1000~r_{\rm{g}}$, and a single emissivity profile ($\propto r^{-q}$, $q$ being the emissivity index) was considered throughout the whole disk with $q$ fixed at the canonical value of 3. The photon index ($\Gamma$) and electron temperature ($kT_{\rm{e}}$) were varied independently of those in thcomp. Other free parameters of relxillCp were the inner disk radius ($R_{\rm{in}}$), log of disk ionization ($\log\xi$), iron abundance ($A_{\rm{Fe}}$), log of disk density ($\log N$), reflection fraction ($R_{\rm{ref}}$), and normalization. To avoid the over prediction of the contribution of the reflection at lower energies, we multiplied relxillCp with a mbknpo component, which helps to put a break on the reflection spectrum below a certain energy, and thus preventing the unphysical runaway of the photon spectrum at low energies (Steiner et al., 2024; Svoboda et al., 2024). The mbknpo component is parameterized with the break energy $B$ and index $I$, used for the correction for energies $E<B$. We set these parameters according to the prescription outlined in Svoboda et al. (2024). This composite model, plabs*tbabs(thcomp*diskbb+mbknpo*relxillCp) (hereafter M1), resulted in a significantly improved fit with $\chi^{2}/dof=457.3/590$. The best-fit spectral parameters are listed in Table 2, and the corresponding spectral models and data are shown in Figure 4.
@@
The results from our model M1 clearly indicate that the broadband spectrum in the 0.7780.7-78 keV band in the HIMS requires two separate Comptonizing regions. The relxillCp component represents the hard Comptonizing region with $\Gamma\sim 1.8$ and $kT_{\rm{e}}\sim 16$ keV, likely having a spherically extended geometry. This region is primarily responsible for the origin of the reflection features at a distant part of the disk. On the other hand, the second Comptonizing region, described by thcomp, is found to have low electron temperature, $kT_{\rm{e}}\sim 4$ keV, and exhibits a relatively steeper spectrum with $\Gamma\sim 1.95$. This soft Comptonizing region may appear as a warm layer over the inner parts of the accretion disk. We also explored the possibility of an additional contribution to the reflection from the soft Comptonizing region by incorporating a second relxillCp component as a reflection-only component. However, we did not find any significant improvement in our spectral fitting. This implies that the reflection from the soft Comptonizing region is negligible and that the hard Comptonizing region dominates the overall reflection spectrum.
@@
The above model could not properly constrain the inner disk radius, $R_{\rm{in}}$. However, the lower bound at the $90\%$ confidence level indicates that the disk is truncated away from the ISCO by at least $\sim 33~r_{\rm{g}}$ (where $r_{\rm{g}}=GM/c^{2}$ is the Gravitational radius). We also calculated the inner disk radius from the best-fit value of the diskbb normalization, following the XSPEC definition and applying the color correction factor of 1.7 (Shimura and Takahara, 1995). The estimated radius is $\sim 29~r_{\rm{g}}$, which agrees closely with the value obtained from the reflection modeling. Furthermore, the temperature of the inner accretion disk is estimated to be $\sim 0.4$ keV, typically higher than that observed in the LHS. The disk appears highly ionized, with slightly elevated density. Within the uncertainties, the iron abundance is consistent with the solar value. Additionally, the reflection fraction, $R_{\rm{ref}}$, defined as the ratio of the coronal emission reaching the disk and the observer, is found to be $\sim 0.64$, which in turn also favors the disk truncation scenario. All the best-fit spectral parameters indicate that the source was in HIMS during Obs1.
@@
We then modeled the hard Comptonization spectra with the compps (Poutanen and Svensson, 1996) model by replacing relxillCp in M1. The Comptonization model compps can treat the electron distribution in a corona as purely thermal, purely non-thermal, or hybrid for different shapes of the coronal geometry. Here, we assumed the electron distribution in the corona as purely thermal and the coronal geometry as spherical (geom=0) in the compps model. We also tested other corona geometries but observed no significant change in the spectral fit. Therefore, we kept the geometry spherical, which also yields a faster method for fitting the spectra than the other geometries. In the compps, the seed photons for the Comptonization are considered to be multi-color disk photons provided by the diskbb component. We fixed the disk inclination angle at $40\arcdeg$ and iron abundance at solar. The other free parameters of this model component are the electron temperature, $kT_{\rm{e}}$, and the optical depth, $\tau$, of the scattering medium. Furthermore, to account for disk reflection, we convolved the xilconv model with both Comptonizing regions. The reflection fraction, $R_{\rm{refl}}$, and log of disk ionization, $\log\xi$, parameters in compps were allowed to vary freely. We kept the iron abundance fixed at solar, consistent with the value obtained from model M1 within the uncertainty. We also incorporated the relconv model to treat the relativistic blurring effects, while keeping the relevant parameters the same as those used for relxillCp in model M1. This model, plabs*tbabs*relconv*xilconv*(thcomp*diskbb+compps) (referred to as M2), provides a statistically acceptable fit to the broadband spectrum, with $\chi^{2}/dof=458.7/592$. The best-fit spectral parameters are given in Table 2, and the best-fit spectra along with model components are shown in Figure 4.
@@
In the next scenario, we incorporated spectral data from the HE and CZTI instruments onboard Insight-HXMT and AstroSat, respectively, to achieve broadband spectral coverage up to 200 keV. We first assessed calibration uncertainties in these datasets by fitting them up to 78 keV, along with NICER and NuSTAR spectra, using both models M1 and M2. The resulting best-fit spectral parameters were either identical or consistent with those obtained from joint spectral fitting of NICER and NuSTAR data in the 0.7–78 keV band using models M1 and M2, indicating no significant cross-calibration discrepancies among the instruments. Furthermore, inspection of the HE spectral data confirms that the background is significantly lower than the source signal in the 100150100-150 keV band, with source and background count rates of $96.7\pm 0.2$c/s and $48.1\pm 0.9$ c/s, respectively. This suggests a detection significance of $\sim 52\sigma$ for the source emission in the 100150100-150 keV band. We then extended the energy ranges of the HE and CZTI spectral data to $150$ keV and $200$ keV, respectively. However, we find that both models, M1 and M2, were unable to describe the broadband spectra up to $200$ keV and yielded poor spectral fits with $\chi^{2}$/dof=803/679=803/679 for M1 and $886.8/682$ for M2, leaving a prominent hump at high energies ($\gtrsim 100$ keV). Figure 5 presents the broadband spectral data in the 0.72000.7-200 keV band fitted with the model M2, along with the residuals, highlighting the deviations between the observed data and the model.